WHY DISTANCE IS CONSIDERED SQUARE IN COULOMB’S LAW


1
Expert's answer
2021-04-28T09:26:52-0400

Answering your question with analogy to Newton's Law of Gravitation,

Consider a charged object moving in a circle with constant speed,

"F=centripetal\space force"

"F=\dfrac{mv^2}{r}"


For an object moving in circle,

"v=\dfrac{2\pi r}{T}"

Substituting the value of "v" in above equation

"F=\dfrac{m(\dfrac{2\pi r}{T})^2}{r}"

"F=\dfrac{4m\pi^2r}{T^2}"

"F=4m\pi^2\dfrac{r}{T^2}"

From Kepler's third law,

"T^2\propto r^3"

"T^2=kr^3"


Substituting the value of "T^2" , in above equation,

"F=4m\pi^2\dfrac{r}{kr^3}"

"F=\dfrac{4m\pi^2}{k}\times\dfrac{1}{r^2}"

"F\propto\dfrac{1}{r^2}"